OpenStudy (anonymous):

How do you integrate from 1 to 2 ((x^2-1)/(x+1)). Please.

OpenStudy (anonymous):

factor x^2 - 1 = (x+1)(x-1) cancel stuff out and your left with this to integrate \[\int\limits_{1}^{2} (x-1) dx\] so the antiderivative is x^2/2 - x and now you substitute the limits in (2)^2/2 - 2 - [(1)^2/2 -1] 0 - (1/2 - 1) = 1/2

OpenStudy (anonymous):

ohhhhhhhhhh..... I knew that (jk) thanks a lot

OpenStudy (anonymous):

np glad to help